# ElementaryNet: A Non-Strategic Neural Network for Predicting Human Behavior in Normal-Form Games

### v2

@@ Abstract
Behavioral game theory models serve two purposes: yielding insights into how human decision-making works, and predicting how people would behave in novel strategic settings. A system called GameNet represents the state of the art for predicting human behavior in the setting of unrepeated simultaneous-move games, combining a simple “level-kk” model of strategic reasoning with a complex neural network model of non-strategic “level-0” behavior. Although this reliance on well-established ideas from cognitive science ought to make GameNet interpretable, the flexibility of its level-0 model raises the possibility that it is able to emulate strategic reasoning. In this work, we prove that GameNet’s level-0 model is indeed too general. We then introduce ElementaryNet, a novel neural network that is provably incapable of expressing strategic behavior. We show that these additional restrictions are empirically harmless, leading ElementaryNet to statistically indistinguishable predictive performance vs GameNet. We then show how it is possible to derive insights about human behavior by varying ElementaryNet’s features and interpreting its parameters, finding evidence of iterative reasoning, learning about the depth of this reasoning process, and showing the value of a rich level-0 specification.
@@ 1 Introduction
Human behavior in strategic settings often deviates significantly from the predictions of classical game theory: for instance, humans often play dominated actions (as in the unrepeated Prisoner’s Dilemma), which a fully rational agent would not. Behavioral game theory aims to address these shortcomings by developing predictive models of human strategic behavior. The purpose of such models is two-fold. First, such models that have been trained on experimental data can be analyzed to yield insight about human psychology. In this sense, classical game theory can be seen as a zero-parameter model that predicts poorly; the goal here is to find models that are as simple and interpretable as possible, and that make much more accurate predictions. Second, we might simply care about building models that are as accurate as possible. Such models are useful for developing agents that will interact with humans or for designing systems of rules (“mechanisms”) that aim for good performance when faced with real human participants.
@@
Past work has made substantial progress towards the first goal, demonstrating that so-called iterative reasoning models with just two parameters are surprisingly good at predicting behavior in new strategic settings, and hence giving insight into the way human subjects reason. Let us describe the best performing such model, Quantal Cognitive Hierarchy (QCH) (Camerer, Ho, and Chong 2004). QCH predicts that people perform strategic reasoning at some finite “level”. The model includes a probability distribution describing the proportion of agents at each level; when this distribution is Poisson (having one parameter), we call the model QCHp. Level-0 reasoners are nonstrategic: they perform some arbitrary computation that falls short of forming beliefs about their opponents and best responding to these beliefs. Camerer, Ho, and Chong (2004) simply assert that level-0 reasoners choose actions uniformly at random, which is clearly nonstrategic (and parameter free); we call this instantiation of the model Uniform + QCHp. Level-1 reasoners quantally best respond to level-0 agents, where quantal response is a noisy version of best response that depends on a “precision” parameter. Level-$k$ reasoners for $k\geq 2$ are more complex: they know the distribution over levels $0,\ldots,k-1$, and quantally best respond to the aggregate distribution of play by all lower-level agents. More recent work showed that even better performance can be achieved by hand-crafting a richer level-0 model based on insights from cognitive psychology, at the expense of adding more parameters to the model (Wright and Leyton-Brown 2019).
@@
Regarding the second goal of raw predictive performance, it is perhaps unsurprising that the current state-of-the-art model for predicting human behavior in unrepeated normal form games, dubbed GameNet (Hartford, Wright, and Leyton-Brown 2016), is based on a neural network. GameNet extends the Uniform + QCHp architecture by replacing the uniform level-0 specification with a learned model and the Poisson distribution by a finite-depth histogram. The learned model uses a permutation-equivariant architecture to express the inductive bias that an action should be played with the same probability regardless of its index in the game matrix, but is otherwise unrestricted. Empirically, GameNet dramatically outperforms Uniform + QCHp. Although its neural network component is an uninterpretable black box, GameNet was based on the QCH architecture in part to allow for meaningful interpretation: one can ask about the shape of the level distribution; the precision parameter; and the way model performance varies when levels are added or removed. In this vein, Hartford, Wright, and Leyton-Brown (2016) made a striking observation: GameNet performed best when it was restricted to predict that all agents were level-0. What should we make of this result? Does it show that iterative strategic reasoning is a dead end, and that human subjects are better described in some different way? Or does it show that a sufficiently general neural network is already able to simulate iterative strategic reasoning, meaning that such a network has nothing to gain from the explicit addition of QCH levels, and hence that GameNet’s apparent interpretability was a mirage? Finally, how would we tell the difference?
@@
In this work, we make three main contributions. First, in Section˜3, we show that GameNet’s purportedly level-0 neural network is capable of strategic reasoning, and is hence inappropriate to use for describing level-0 agents. Our proof is constructive: we give a specific setting of its parameters that computes quantal best response to maxmax, a strategic model. This finding helps to explain why removing levels of strategic reasoning did not improve GameNet’s performance.
@@
Second, in Section˜4, we introduce ElementaryNet, a new restriction of GameNet’s architecture that is only capable of non-strategic behavior. Our architecture is rooted in the concept of elementary models (Wright and Leyton-Brown 2022), a set of behavioral models which are provably incapable of representing strategic behavior. Elementary models compute a real-valued “potential” for each outcome of the game that summarizes the set of utilities for all the players into a single number, and then compute a distribution of play using only those summaries. Intuitively, this restriction to a single summary per outcome prevents the model from forming a belief about the opponent’s play (by assuming that the opponent will optimize her own utilities), and then best responding to that belief (by optimizing the model’s own utilities). Our architecture is a convex combination of neural networks whose inputs are summarized by a single potential function each; it is thus a convex combination of elementary models, which is provably non-strategic in a specific formal sense.
@@
Finally, in Section˜5, we perform extensive experiments on ElementaryNet. First, we show that, despite the additional restrictions on its architecture, ElementaryNet combined with an iterative reasoning model achieves predictive power statistically indistinguishable from that of GameNet. Second, we show how the model’s clean delineation between strategic and non-strategic components makes it possible to obtain interpretable insights about human behavior by varying different features of the model. In particular, we show that ElementaryNet performs significantly worse without a strategic model, demonstrating that iterative reasoning is indeed a good model of human behavior. We show that restricting the level-0 model to only consider the player’s own payoffs degrades performance, demonstrating the value of a rich level-0 specification. Conversely, we also show that allowing the model to learn the exact potentials that summarize player utilities does not improve performance when compared to a natural set of four basis potentials. Finally, we successfully interpret the parameters of the iterative reasoning model that are co-learned with ElementaryNet, in exactly the way that was uninformative with GameNet.
@@ 2 Preliminaries
A behavior is an element of the simplex $s_{i}\in\Delta(A_{i})$, representing a distribution over agent $i$’s actions; we use the non-standard term “behavior”, rather than the more standard “strategy”, to avoid other awkward terminology, such as a “non-strategic strategy”. A behavior profile is a tuple of behaviors $s=(s_{1},s_{2})$. Overloading notation, we denote an agent’s expected utility as $u_{i}(s)$ = $\mathds{E}_{a\sim s}u_{i}(a)$. For either agent $i$, we write $s_{-i}$ to represent the behavior of the other agent, and $(s_{i},s_{-i})$ to refer to a behavior profile. A behavioral model is a function $f_{i}$ that maps a game $G=(A,u)$ to a behavior $f_{i}(G)\in\Delta(A_{i})$.
@@
We now describe several behavioral models used in prior behavioral game theory work. A common building block in many of these models is the concept of quantal best-response. {definition} The (logit) quantal best-response to a strategy $s_{-i}$ is $QBR_{i}(s_{-i};\lambda,G)(a_{i})\propto\exp[\lambda\cdot u_{i}(a_{i},s_{-i})],$ where $\lambda$ (the precision parameter) controls the agent’s sensitivity to differences in utilities.
The quantal cognitive hierarchy (QCH) model (e.g., Wright and Leyton-Brown 2017) combines quantal best response with a model of iterative reasoning. {definition} Let $G$ be a game, $\lambda\in\mathds{R}$ be a precision, $s^{0}$ be a profile of level-0 behaviors in $G$, and $D$ be a probability distribution over levels. Then, the level-$k$ quantal hierarchical behavior is defined as $s^{k}_{i}=QBR_{i}(s_{-i}^{0:k-1};\lambda,G),$ where $s^{0:k-1}_{i}(a_{i})\propto\sum_{m=0}^{k-1}D(m)s^{m}_{i}(a_{i})$. The quantal cognitive hierarchy behavior is the weighted average of these behaviors $QCH_{i}(a_{i})=\sum_{k}D(k)s^{k}_{i}(a_{i})$.
QCH models depend critically on the level-0 model, which determines not just the behavior of level-0 agents, but also of higher-level agents, who react to it. A natural, simple choice is the uniform distribution, which was originally used in the Level-$k$ (Nagel 1995; Costa-Gomes, Crawford, and Broseta 2001) and Cognitive Hierarchy (Camerer, Ho, and Chong 2004) models. However, Wright and Leyton-Brown (2019) found improvements using richer level-0 models based on heuristics from cognitive science. For example, one such heuristic is the maxmax behavioral model,
@@
Notably, this heuristic depends only on the agent’s utility $u_{i}$, neglecting the opponent’s utility $u_{-i}$. The other heuristics (described in the appendix) also depend on simple linear combinations of the players’ utilities, such as the sum (or welfare) $u_{1}+u_{2}$ or the difference (unfairness) $u_{1}-u_{2}$.
@@
GameNet (Hartford, Wright, and Leyton-Brown 2016) is a deep learning architecture for predicting human strategic behavior. It can be understood as taking to the logical limit the insight that a richer level-0 specification can yield better performance. Broadly, GameNet consists of two parts: feature layers, which are intended to model level-0 behavior, and action response (AR) layers, which perform iterative strategic reasoning. AR layers may be understood as a generalization of the QCH model, with additional parameters independently controlling each level’s precision, the distribution over lower levels to which agents respond, and the transformed utility matrices which they use for this response. We devote more attention to the feature layers.
@@
After the final hidden layer, the matrices $\{H^{L,c}\}_{c=1:C_{L}}$ are transformed into a single distribution over the row player’s actions as $f=\sum_{c=1}^{C_{L}}w_{c}f^{c}$, where $f^{c}_{i}=\operatorname*{softmax}(\sum_{j}H^{L,c}_{i,j})$, $\operatorname*{softmax}(x)_{i}=\exp(x_{i})/\sum_{i^{\prime}}\exp(x_{i^{\prime}})$, and the weights $w_{c}\in\Delta(C_{L})$ are subject to simplex constraints. An analogous predicted distribution over the column player’s actions is made by replacing the input utility matrices $U^{1}$ and $U^{2}$ with $(U^{2})^{T}$ and $(U^{1})^{T}$, respectively. This architecture is summarized in Figure˜1.
Compared to a more standard, off-the-shelf architecture—e.g., flattening the utility matrices into a vector of length $2nm$ and applying a feedforward neural network—feature layers have several advantages. One is that they are permutation equivariant, which guarantees that permuting the utility matrices will permute the predictions correspondingly. Permutation equivariance lowers the number of parameters of the network, making it learn more efficiently and removing the need for data augmentation. Feature layers are also agnostic to the size of the game, making it possible to learn from heterogeneous data with a variety of action spaces or generalize to games of new sizes.
@@
To make precise claims about whether or not a model is strategic, we adopt a formal definition of (non-)strategic behavior from Wright and Leyton-Brown (2022). In particular, we use their definition of a weakly non-strategic model. {definition} We say that an action $a_{i}^{+}$ in a game $G=(A,u)$ is $\zeta$-dominant if $u_{i}(a_{i}^{+},a_{-i})>u_{i}(a_{i},a_{-i})+\zeta$ for all $a_{i}\neq a_{i}^{+}$ and $a_{-i}\in A_{-i}$. Then, a behavioral model $f_{i}$ is dominance responsive if there exists some $\zeta>0$ such that, in all games $G$ with a $\zeta$-dominant action $a_{i}^{+}$, the mode of $f_{i}$ is $a_{i}^{+}$: that is, $f_{i}(G)(a_{i}^{+})>f_{i}(G)(a_{i})$ for all $a_{i}\neq a_{i}^{+}$. {definition} A behavioral model $f_{i}$ is weakly non-strategic if it cannot be represented as quantal best response to some dominance-responsive model $f_{-i}$.
Wright and Leyton-Brown (2022) also define a large class of non-strategic behavioral models called elementary models. These models independently compute a “potential” for each outcome in the game, then predict an action distribution based only on these potentials. Intuitively, compressing two utilities into a single potential creates an information bottleneck, which stops the model from representing strategic behavior. Formalizing this intuition requires the following two definitions, which make it possible to ensure that the input utilities cannot be reliably reconstructed from the potentials. {definition} A function $\varphi:\mathds{R}^{2}\to\mathds{R}$ is dictatorial iff it is completely determined by one input: that is, either $\varphi(x,y)=\varphi(x,y^{\prime})$ for all $x,y,y^{\prime}\in\mathds{R}$, or $\varphi(x,y)=\varphi(x^{\prime},y)$ for all $x,x^{\prime},y\in\mathds{R}$. {definition} A function $\varphi:\mathds{R}^{2}\to\mathds{R}$ is non-encoding iff, for all $i\in\{1,2\}$ and $b>0$, there exist $x,x^{\prime}\in\mathds{R}^{2}$ such that $\varphi(x)=\varphi(x^{\prime})$ but $|x_{i}-x^{\prime}_{i}|>b$. {definition} An elementary behavioral model is a model of the form $f_{i}(G)=h_{i}(\Phi(G))$, where
@@
The key property of elementary models that we will leverage for our results is that convex combinations of elementary models are weakly non-strategic.
@@
(Wright and Leyton-Brown 2022, Theorem 5.) Let $f_{i}(G)=\sum_{k=1}^{K}w_{k}g_{i}^{k}(G)$ be a convex combination of elementary behavioral models. Then, $f_{i}$ is weakly non-strategic.

### v3

@@ Abstract
Behavioral game theory models serve two purposes: yielding insights into how human decision-making works, and predicting how people would behave in novel strategic settings. A system called GameNet represents the state of the art for predicting human behavior in the setting of unrepeated simultaneous-move games, combining a simple “level-kk” model of strategic reasoning with a complex neural network model of non-strategic “level-0” behavior. Although this reliance on well-established ideas from cognitive science ought to make GameNet interpretable, the flexibility of its level-0 model raises the possibility that it is able to emulate strategic reasoning. In this work, we prove that GameNet’s level-0 model is indeed too general. We then introduce ElementaryNet, a novel neural network that is provably incapable of expressing strategic behavior. We show that these additional restrictions are empirically harmless, with ElementaryNet and GameNet having statistically indistinguishable performance. We then show how it is possible to derive insights about human behavior by varying ElementaryNet’s features and interpreting its parameters, finding evidence of iterative reasoning, learning about the depth of this reasoning process, and showing the value of a rich level-0 specification.
Code & datahttps://github.com/gregdeon/elementarynet
@@ 1 Introduction
Human behavior in strategic settings often deviates significantly from the predictions of classical game theory: for instance, humans often play dominated actions (as in the unrepeated Prisoner’s Dilemma), which a fully rational agent would not. Behavioral game theory aims to address these shortcomings by developing predictive models of human strategic behavior. The purpose of such models is two-fold. First, such models that have been trained on experimental data can be analyzed to yield insight about human psychology. In this sense, classical game theory can be seen as a zero-parameter model that predicts poorly; the goal here is to find models that make much more accurate predictions, but are as simple and interpretable as possible. Second, we might simply build models that are as accurate as possible. Such models are useful for developing agents that will interact with humans or for designing systems of rules (“mechanisms”) that will perform well when faced with human participants.
@@
Past work has made substantial progress towards the first goal, demonstrating that so-called iterative reasoning models with just two parameters are surprisingly good at predicting behavior in new strategic settings, and hence giving insight into the way human subjects reason. Let us describe Quantal Cognitive Hierarchy (QCH), the best performing such model (Camerer2004). QCH predicts that people perform strategic reasoning at some finite “level”. The model includes a probability distribution describing the proportion of agents at each level; when this distribution is Poisson (having one parameter), we call the model QCHp. Level-0 reasoners are nonstrategic: they perform some arbitrary computation that falls short of forming beliefs about their opponents and best responding to these beliefs. Camerer2004 simply assert that level-0 reasoners choose actions uniformly at random, which is clearly nonstrategic (and parameter free); we call this instantiation of the model Uniform + QCHp. Level-1 reasoners quantally best respond to level-0 agents, where quantal response is a noisy version of best response that depends on a “precision” parameter. Level-$k$ reasoners for $k\geq 2$ are more complex: they know the distribution over levels $0,\ldots,k-1$, and quantally best respond to the aggregate distribution of play by all lower-level agents. More recent work showed that even better performance can be achieved by hand-crafting a richer level-0 model based on insights from cognitive psychology, at the expense of adding more parameters (JAIR).
@@
Regarding the second goal of raw predictive performance, it is perhaps unsurprising that the current state-of-the-art model for predicting human behavior in unrepeated normal-form games, dubbed GameNet (GameNet), is based on a neural network. GameNet extends the Uniform + QCHp architecture by replacing the uniform level-0 specification with a learned model and the Poisson distribution by a finite-depth histogram. The learned model uses a permutation-equivariant architecture to express the inductive bias that an action should be played with the same probability regardless of its index in the game matrix, but is otherwise unrestricted. Empirically, GameNet dramatically outperforms Uniform + QCHp.
Although its neural network component is an uninterpretable black box, GameNet was based on the QCH architecture in part to allow for meaningful interpretation: one can ask about the shape of the level distribution, the precision parameter, and the way model performance varies when levels are added or removed. In this vein, GameNet made a striking observation: GameNet performed best when it was restricted to predict that all agents were level-0. What should we make of this result? Does it show that iterative strategic reasoning is a dead end, and that human subjects are better described in some different way? Or, does it show that a sufficiently general neural network is already able to simulate iterative strategic reasoning, and hence that GameNet’s apparent interpretability was a mirage? Finally, how would we tell the difference?
@@
In this work, we make three main contributions. First, in Section˜3, we show that GameNet’s purportedly level-0 neural network is capable of strategic reasoning, and is hence inappropriate to use for describing level-0 agents. Our proof is constructive: we give a specific setting of its parameters that computes quantal best response to maxmax, a strategic model. This finding helps to explain why adding levels of strategic reasoning did not improve GameNet’s performance.
@@
Second, in Section˜4, we introduce ElementaryNet, a new restriction of GameNet’s architecture that is only capable of non-strategic behavior. Our architecture is rooted in the concept of elementary models (WLBnonstrategic), a set of behavioral models which are provably incapable of representing strategic behavior. Elementary models compute a real-valued “potential” for each outcome of the game that summarizes the set of utilities for all the players into a single number, and then compute a distribution of play using only those potentials. Intuitively, this restriction to a single summary per outcome prevents the model from forming a belief about the opponent’s play (based on the opponent’s utilities), and then best responding to that belief (based on the model’s own utilities). Our architecture is a convex combination of neural networks whose inputs are summarized by a single potential function each; it is thus a convex combination of elementary models, which is provably non-strategic in a precise, formal sense.
@@
Finally, in Section˜5, we perform extensive experiments on ElementaryNet. We first show that, despite the additional restrictions on its architecture, ElementaryNet combined with an iterative reasoning model achieves predictive power statistically indistinguishable from that of GameNet. We then show how the model’s clean delineation between strategic and non-strategic components makes it possible to obtain interpretable insights about human behavior by varying different features of the model. In particular, we show that ElementaryNet performs significantly worse without a strategic model, demonstrating that iterative reasoning is indeed a good model of human behavior. We show that restricting the level-0 model—to only consider the player’s own payoffs, or to use a fixed set of four basis potentials—degrades performance, demonstrating the value of a rich level-0 specification. Finally, we successfully interpret the parameters of the iterative reasoning model that are co-learned with ElementaryNet, in exactly the same way that was uninformative with GameNet.
@@ 2 Preliminaries
A behavior $s_{i}$ is an element of the simplex $\Delta(A_{i})$, representing a distribution over agent $i$’s actions; we use the non-standard term “behavior”, rather than the more standard “strategy”, to avoid other awkward terminology, such as a “non-strategic strategy”. A behavior profile is a tuple of behaviors $s=(s_{1},s_{2})$. Overloading notation, we denote an agent’s expected utility as $u_{i}(s)$ = $\mathds{E}_{a\sim s}u_{i}(a)$. For either agent $i$, we write $s_{-i}$ to represent the behavior of the other agent, and $(s_{i},s_{-i})$ to refer to a behavior profile. A behavioral model is a function $f_{i}$ that maps a game $G=(A,u)$ to a behavior $f_{i}(G)\in\Delta(A_{i})$.
@@
We now describe several behavioral models used in prior behavioral game theory work. A common building block in many of these models is the concept of quantal best response. {definition} The (logit) quantal best response to a strategy $s_{-i}$ is $QBR_{i}(s_{-i};\lambda,G)(a_{i})\propto\exp[\lambda\cdot u_{i}(a_{i},s_{-i})],$ where $\lambda$ (the precision parameter) controls the agent’s sensitivity to differences in utilities.
The quantal cognitive hierarchy (QCH) model (e.g., WLB2017) combines quantal best response with a model of iterative reasoning. {definition} Let $G$ be a game, $\lambda\in\mathds{R}$ be a precision, $s^{0}$ be a profile of level-0 behaviors in $G$, and $D$ be a probability distribution over levels. Then, the level-$k$ quantal hierarchical behavior is defined as $s^{k}_{i}=QBR_{i}(s_{-i}^{0:k-1};\lambda,G),$ where $s^{0:k-1}_{i}(a_{i})\propto\sum_{m=0}^{k-1}D(m)s^{m}_{i}(a_{i})$. The quantal cognitive hierarchy behavior is the weighted average of these behaviors $QCH_{i}(a_{i})=\sum_{k}D(k)s^{k}_{i}(a_{i})$.
QCH models depend critically on the level-0 model, which determines not just the behavior of level-0 agents, but also of higher-level agents who react to it. A natural, simple choice is the uniform distribution, which was originally used in the Level-$k$ (Nagel1995; CostaGomes2001) and Cognitive Hierarchy (Camerer2004) models. However, JAIR found improvements using richer level-0 models based on heuristics from cognitive science. For example, one such heuristic is the maxmax behavioral model,
@@
Notably, this heuristic depends only on the agent’s utility $u_{i}$, neglecting the opponent’s utility $u_{-i}$. The other heuristics (see Appendix A) also depend on simple linear combinations of the players’ utilities, such as the sum (or welfare) $u_{1}+u_{2}$ or the difference (unfairness) $u_{1}-u_{2}$.
@@
GameNet (GameNet) is a deep learning architecture for predicting human strategic behavior. It can be understood as taking the insight that a richer level-0 specification can yield better performance to its logical limit. Broadly, GameNet consists of two parts: feature layers, which are intended to model level-0 behavior, and action response (AR) layers, which perform iterative strategic reasoning. AR layers are essentially a generalization of the QCH model, with additional parameters independently controlling each level’s precision, the distribution over lower levels to which agents respond, and the transformed utility matrices which they use for this response. We devote more attention to the feature layers.
@@
After the final hidden layer, the matrices $\{H^{L,c}\}_{c=1:C_{L}}$ are transformed into a single distribution over the row player’s actions as $f=\sum_{c=1}^{C_{L}}w_{c}f^{c}$, where $f^{c}=\operatorname*{softmax}(\sum_{j}H^{L,c}_{i,j})$, $\operatorname*{softmax}(x)_{i}=\exp(x_{i})/\sum_{i^{\prime}}\exp(x_{i^{\prime}})$, and the weights $w_{c}\in\Delta(C_{L})$ are subject to simplex constraints. An analogous predicted distribution over the column player’s actions is made by replacing the input utility matrices $U^{1}$ and $U^{2}$ with $(U^{2})^{T}$ and $(U^{1})^{T}$, respectively. This architecture is summarized in Figure˜1.
Compared to a more standard, off-the-shelf architecture—e.g., flattening the utility matrices into a vector of length $2nm$ and applying a feedforward neural network—feature layers have several advantages. One is that they are permutation equivariant, which guarantees that permuting the utility matrices will permute the predictions correspondingly. Permutation equivariance lowers the number of parameters of the network, making it learn more efficiently and removing the need for data augmentation. Feature layers are also agnostic to the size of the game, making it possible to learn from heterogeneous data with a variety of action spaces and to generalize to games of new sizes.
@@
To make precise claims about whether or not a model is strategic, we adopt a formal definition of (non-)strategic behavior from WLBnonstrategic. In particular, we use their definition of a weakly non-strategic model. {definition} An action $a_{i}^{+}$ in a game $G=(A,u)$ is $\zeta$-dominant if $u_{i}(a_{i}^{+},a_{-i})>u_{i}(a_{i},a_{-i})+\zeta$ for all $a_{i}\neq a_{i}^{+}$ and $a_{-i}\in A_{-i}$. Then, a behavioral model $f_{i}$ is dominance responsive if there exists some $\zeta>0$ such that, in all games $G$ with a $\zeta$-dominant action $a_{i}^{+}$, the mode of $f_{i}$ is $a_{i}^{+}$: that is, $f_{i}(G)(a_{i}^{+})>f_{i}(G)(a_{i})$ for all $a_{i}\neq a_{i}^{+}$. {definition} A behavioral model $f_{i}$ is weakly non-strategic if it cannot be represented as quantal best response to some dominance-responsive model $f_{-i}$.
WLBnonstrategic also define a class of non-strategic behavioral models called elementary models. These models independently compute a “potential” for each outcome in the game, then predict an action distribution based only on these potentials. Intuitively, as long as the potential function is well-behaved, compressing two utilities into one potential creates an information bottleneck that prevents the model from representing strategic behavior. {definition} A function $\varphi:\mathds{R}^{2}\to\mathds{R}$ is dictatorial iff it is completely determined by one input: that is, either $\varphi(x,y)=\varphi(x,y^{\prime})$ for all $x,y,y^{\prime}\in\mathds{R}$, or $\varphi(x,y)=\varphi(x^{\prime},y)$ for all $x,x^{\prime},y\in\mathds{R}$. {definition} A function $\varphi:\mathds{R}^{2}\to\mathds{R}$ is non-encoding iff, for all $i\in\{1,2\}$ and $b>0$, there exist $x,x^{\prime}\in\mathds{R}^{2}$ such that $\varphi(x)=\varphi(x^{\prime})$ but $|x_{i}-x^{\prime}_{i}|>b$. {definition} An elementary behavioral model is a model of the form $f_{i}(G)=h_{i}(\Phi(G))$, where
@@
The key fact that we will leverage is that convex combinations of elementary models are weakly non-strategic.
@@
(Theorem 5, WLBnonstrategic) Let $g_{i}^{1},\dots,g_{i}^{K}$ be elementary behavioral models, and let $w_{1},\dots,w_{K}$ be non-negative weights summing to 1. Then, the convex combination $f_{i}(G)=\sum_{k=1}^{K}w_{k}g_{i}^{k}(G)$ is a weakly non-strategic behavioral model.